OpenStudy (anonymous):

On a standardized test with a normal distribution, the mean was 64.3 and the standard deviation was 5.4. what percentile is a 74?

OpenStudy (anonymous):

You'll need to use a Z-table for this one. First, Z = (X - mean)/(Standard deviation), where X is the given value. Z = (74-64.3)/5.4= 1.80 Next, look up 1.80 in the Z table.

OpenStudy (anonymous):

The table value at Z = 1.80 is 0.0359. The percentile can be calculated as (1-0.0359)*100=96.41 percentile

OpenStudy (anonymous):

ok next between what two values does 68.2% of the distribution lie?

OpenStudy (anonymous):

68.2% of data will fall within one standard deviation of the mean

OpenStudy (anonymous):

so, 64.3-5.4=58.9 is the lower bound and 64.3+5.4=69.7 is the upper

OpenStudy (anonymous):

ok thanks